# Efficient Monte-Carlo sampling of metastable systems using non-local collective variable updates

### v2

@@ I Introduction
Enhanced sampling algorithms go beyond the naive Markov Chain Monte Carlo approach to tackle metastability, see Ref. Henin_Lelievre_Shirts_Valsson_Delemotte_2022 for a review in the context of molecular dynamics. A first class of methods originated from the observation that the metastability is removed at sufficiently high temperatures, which led to the development of algorithms like parallel tempering, annealed importance sampling, and sequential Monte Carlo swendsenReplicaMonteCarlo1986; nealAnnealedImportanceSampling2001; delmoralSequentialMonteCarlo2006; woodard_conditions_2009; syed_non-reversible_2022. While known to work on a wide range of problems, they spend a substantial amount of computation time at temperatures not of direct interest and generally require a careful choice of parameters (e.g. for parallel tempering the temperature ladder and the frequency of exchange attempts).
@@
Another important category of algorithms makes use of a collective variable (CV), a lower-dimensional representation of the system that is supposed to capture the metastability of the system. A ‘good’ CV that resolves the metastability can be used for biasing simulations to bridge modes, using enhanced sampling algorithms such as umbrella samplingtorrie-valleau-77 and related histogram methodsKBSKR92; SC08, thermodynamic integrationkirkwood-35 or free energy adaptive biasing methodsdarve-pohorille-01; henin-chipot-04; laio-parrinello-02; barducci-bussi-parrinello-08. However, identifying a suitable CV in low dimension is by no means trivial and thus poses the main limitation to this class of methods.
@@
In recent years, another approach has emerged, which is exploiting generative machine learning models. Among the various generative models, normalizing flowspapamakarios2021 are particularly well-suited for the purposes of MCMC sampling since their tractable probability density allows to employ them as a non-local proposal density of physical configurations, where any imperfections in training are corrected for via a Metropolis-Hastings accept/reject scheme or self-normalized importance samplingnoeBoltzmannGeneratorsSampling2019; greniouxSamplingApproximateTransport2023b. Thanks to the impressive flexibility of neural networks, proofs of concept have been made in different domains ranging from molecular systems to field theories in quantum chromodynamicsliNeuralNetworkRenormalization2018; noeBoltzmannGeneratorsSampling2019; gabrieAdaptiveMonteCarlo2022; invernizzi_skipping_2022; abbott_sampling_2022. However, the limiting factor to scale these approaches appears to be the ability of the generative model to be precise enough either when the dimension of the problem becomes of the order of a few thousands or when the distribution has singularitiesdeldebbioEfficientModellingTrivializing2021; greniouxSamplingApproximateTransport2023b; schonleOptimizingMarkovChain2023; abbottAspectsScalingScalability2022.Other recent approaches make use for example of transformers bera_accurate_2025 and diffusion models lewisScalableEmulationProtein2025b; bera_how_2025, but these do not give easy access to the underlying probability density and are therefore less suitable for MCMC sampling than normalizing flows.
@@
A promising approach is therefore to combine the strength of CV-enhanced sampling and generative models (normalizing flows) with a CV of ‘intermediate’ dimension (from a dozen to a few hundred degrees of freedom). Within this dimensionality, identifying a CV such that it resolves the metastability will be much easier than finding a one to three-dimensional representation required by most classical CV-based approaches. Conversely, a normalizing flow in an ‘intermediate’ dimension can be trained to much higher accuracy than if one attempted to learn the distribution of the entire system. Since the overall goal is still sampling full configurations of the system, the question arises how to employ such a proposal sampler in CV space to obtain an unbiased sampler for the full distribution. This has been addressed in multiple worksathenesComputationChemicalPotential2002; nilmeier2011nonequilibrium; chen_enhanced_2015; chen2015generalized; neal2005taking, in particular under the names of “Hybrid Nonequilibrium Molecular Dynamics” (HNMD)chen_enhanced_2015; chen2015generalized and “Nonequilibrium candidate Monte Carlo” (NCMC) nilmeier2011nonequilibrium in the context of molecular dynamics, but also in a Bayesian settingkaragiannis_annealed_2013. The central idea is always to drive the system from one value of the CV to another, a procedure referred to as steering in the following, and then accept or reject the resulting proposal in the full-space. A similar procedure has also been used in the simulation of glassy system with shear deformationsdas_annealing_2022. In a previous articleschonleOptimizingMarkovChain2023, we described such an algorithm with driven dynamics in continuous time and identified various time-discretizations for which we proved reversibility with respect to the equilibrium measure. Notably, we also related this to the Jarzynski and the Jarzynski–Crooks equalities jarzynski-97; Crooks99, clarifying the algorithm design. In a parallel worktamagnoneCoarseGrainedMolecularDynamics2024, one version of these algorithms was applied to a toy molecular system, using a normalizing flow as the CV sampler and including an adaptive protocol to train it alongside the MCMC procedure.
@@
The main part of the article divides into two sections: We explain all the theoretical aspects of the problem and the algorithm in Section˜II and go on to demonstrate its properties through several numerical examples in Section˜III. Within Section˜II, the sampling problem is introduced in Section˜II.1, followed by a detailed explanation of the algorithm in Section˜II.2. We state our main theoretical result on the reversibility of the algorithm in Section˜II.3 and conclude in Section˜II.4 with a normalized version of the algorithm and comments on the parameterization. For the numerical studies in Section˜III, we present results on four different model systems of increasing complexity in Sections˜III.1, III.2, III.3 and III.4. We conclude the article and give an outlook on future directions in Section˜IV
@@ II.1 Target distribution and collective variable
$\nu(\mathrm{d}q)=Z_{q}^{-1}e^{-\beta V(q)}\mathrm{d}q$ (II.1)
over positions $q\in\mathbb{R}^{d}$ of dimension $d$, with the energy function $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the inverse temperature $\beta=(k_{B}T)^{-1}>0$ and the normalizing constant $Z_{q}$. Let us assume that a collective variable (CV) $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\ell}$ with $\ell<d$ has been identified and we have a ‘good’ sampler $\rho$ to propose moves in CV space. As mentioned in the introduction, the question of how to use these two ingredients to construct an unbiased MCMC sampler in the full space $\mathbb{R}^{d}$ has previously been addressed in literatureathenesComputationChemicalPotential2002; nilmeier2011nonequilibrium; chen_enhanced_2015; chen2015generalized; neal2005taking, but it is not obvious how to employ these schemes if the map $\xi$ is non-linear. Notably, the geometry of the level sets $\Sigma(z)$ defined as
@@
must be taken into consideration. We present an algorithm which explicitly treats the case of a non-linear CV in the following section. Much of the notation throughout this paper is based on Refs. lelievre-rousset-stoltz-book-10; lelievre_langevin_2012 and we refer the reader to those for an in-depth introduction on the mathematical concepts used here.
@@
$\mu(\mathrm{d}q\,\mathrm{d}p)=Z^{-1}e^{-\beta H(q,p)}\mathrm{d}q\,\mathrm{d}p$ (II.3)
where $Z$ is the normalizing constant, $H(q,p)=\frac{1}{2}p^{\top}M^{-1}p+V(q)$} is the Hamiltonian, and $M\in\mathbb{R}^{d\times d}$ is a symmetric positive definite mass matrix.
@@ II.2.1 Overall structure
The algorithm consists in iterating over $n\in\mathbb{N}$ the following steps: Starting from a state $Q_{n}$ with collective variable value $Z_{n}:=\xi(Q_{n})$, generate $Q_{n+1}$ as follows:
@@ 2. Step 2:
    Generate a new $\widetilde{Q}_{n+1}\in\mathbb{R}^{d}$ that satisfies $\xi(\widetilde{Q}_{n+1})=\widetilde{Z}_{n+1}$ by steering the system in CV space. To this end, (i) specify a schedule to constrain the positions and velocities in CV space, (ii) initialize the position $q^{0}=Q_{n}$ and sample an initial momentum $p^{0}$ and (iii) use this to follow underdamped Langevin dynamics in the RATTLE discretization (all of this to be specified below). These dynamics yield a final state $\widetilde{Q}_{n+1}$ and a quantity $\mathcal{W}_{n+1}$ that is identified as a work.
@@ 3. Step 3:
We explain the various ingredients required for Step 2 in the next three subsections. The idea to sample a system by following some dynamics followed by an accept/reject step that is governed by a work is closely related to the Jarzynski and the Jarzynski–Crooks equalities jarzynski-97; Crooks99.
Remark II.1.
In all of the numerical examples in this work, we will actually use an independent sampler in the CV space, i.e. $\rho(z,z^{\prime})$ is only a function of $z^{\prime}$.
@@ II.2.2 Schedule
To specify the constrained dynamics under which the system evolves from the value of the CV at the previous iteration $Z_{n}$ to the newly proposed value $\widetilde{Z}_{n+1}$, choose first the number of intermediate steps $K_{T}$ associated with this jump. With the fixed timestep $\Delta t$, this sets the physical transition time $T=K_{T}\Delta t$. In practice, we will choose $K_{T}$ as a function of the Euclidean distance $\Delta z=|\widetilde{Z}_{n+1}-Z_{n}|$ of the proposed jump.
@@
for all $k\in\{0,\ldots,K_{T}\}$. Note that the requirement (II.8) precludes the simplest choice for the velocity of $v_{z}(t_{k})=(z(t_{k+1})-z(t_{k}))/\Delta t$. A simple choice is instead a schedule that connects the endpoints on a straight line with a continuously differentiable schedule function $f:[0,1]\rightarrow[0,1]$:
@@
$v^{(Z,\widetilde{Z})}_{z}(t_{k})$ $=(\widetilde{Z}-Z)\frac{\dot{f}(k/K_{T})}{K_{T}\Delta t}.$ (II.10)
The conditions on the schedule translate to conditions on $f$ as $f(0)=0$, $f(1)=1$ and $\dot{f}(0)=\dot{f}(1)=0$, which are fulfilled for example by
@@
See Remark˜B.4 in Section˜B.2.1 for further comments on the velocity schedule. The assumptions (II.6) and (II.8) play a crucial role in the proof of the reversibility of the Markov chain $(Q_{n})_{n\geq 0}$ and we demonstrate in Section˜H.2 how their violation can introduce a bias.
Remark II.2.
@@ II.2.3 Initialization of the momentum
For a given position $q$ and momentum $p$, the velocity in $\xi$-direction is defined as
@@
where $\nabla\xi(q)=\left(\nabla\xi_{1}(q),\ldots,\nabla\xi_{\ell}(q)\right)\in\mathbb{R}^{d\times\ell}$. The initial momentum that is sampled in Step 2 of the algorithm needs to have zero velocity $v_{\xi}$ to agree with condition (II.6) of the velocity schedule. The cotangent space of zero velocity associated with a position $q\in\Sigma(z)$ is
@@
We initialize the momentum $p^{0}$ by sampling an unconstrained momentum $\widetilde{p}^{0}$ from a centered normal distribution with covariance matrix $\frac{1}{\beta}M$ and projecting it onto the required subspace:
@@
II.2.4 RATTLE dynamics and work
To relax the entire configuration $q$ along the chosen CV and velocity schedules, for reasons related to the geometry of the level sets $\Sigma(z)$ (see also the reversibility proof in Section˜B.3), we will follow constrained dynamics with a modified potential energy $\widetilde{V}(q)=V(q)+V_{\mathrm{fix}}(q)$ with the so-called Fixman termFixman1978
$V_{\mathrm{fix}}(q)=\frac{1}{2\beta}\log\left(\det G_{M}(q)\right).$ (II.16)
Starting from $(q^{0},p^{0})$, a discretized path $\left((q^{k},p^{k})_{0\leq k\leq K_{T}}\right)$ is generated by iterating on $0\leq k\leq K_{T}-1$ a version of the RATTLE dynamics with constraints found in literaturelelievre_langevin_2012 (for the sake of readability, we drop the superscript indicating the endpoints of the schedules $z(t)$ and $v_{z}(t)$ in this subsection). The scheme writes as follows:

### v3

@@ I Introduction
Enhanced sampling algorithms go beyond the naive Markov Chain Monte Carlo approach to tackle metastability, see Ref. Hénin et al., 2022 for a review in the context of molecular dynamics. A first class of methods originated from the observation that the metastability is removed at sufficiently high temperatures, which led to the development of algorithms like parallel tempering, annealed importance sampling, and sequential Monte Carlo Swendsen and Wang (1986); Neal (2001); Del Moral et al. (2006); Woodard et al. (2009); Syed et al. (2022). While known to work on a wide range of problems, they spend a substantial amount of computation time at temperatures not of direct interest and generally require a careful choice of parameters (e.g. for parallel tempering the temperature ladder and the frequency of exchange attempts).
@@
Another important category of algorithms makes use of a collective variable (CV), a lower-dimensional representation of the system that is supposed to capture the metastability of the system. A ‘good’ CV that resolves the metastability can be used for biasing simulations to bridge modes, using enhanced sampling algorithms such as umbrella samplingTorrie and Valleau (1977) and related histogram methodsKumar et al. (1992); Shirts and Chodera (2008), thermodynamic integrationKirkwood (1935) or free energy adaptive biasing methodsDarve and Porohille (2001); Hénin and Chipot (2004); Laio and Parrinello (2002); Barducci et al. (2008). However, identifying a suitable CV in low dimension is by no means trivial and thus poses the main limitation to this class of methods.
@@
In recent years, another approach has emerged, which is exploiting generative machine learning models. Among the various generative models, normalizing flowsPapamakarios et al. (2021) are particularly well-suited for the purposes of MCMC sampling since their tractable probability density allows to employ them as a non-local proposal density of physical configurations, where any imperfections in training are corrected for via a Metropolis-Hastings accept/reject scheme or self-normalized importance samplingNoé et al. (2019); Grenioux et al. (2023). Thanks to the impressive flexibility of neural networks, proofs of concept have been made in different domains ranging from molecular systems to field theories in quantum chromodynamicsLi and Wang (2018); Noé et al. (2019); Gabrié et al. (2022); Invernizzi et al. (2022); Abbott et al. (2022a). However, the limiting factor to scale these approaches appears to be the ability of the generative model to be precise enough either when the dimension of the problem becomes of the order of a few thousands or when the distribution has singularitiesDel Debbio et al. (2021); Grenioux et al. (2023); Schönle and Gabrié (2023); Abbott et al. (2022b).Other recent approaches make use for example of transformers Bera and Mondal (2025a) and diffusion models Lewis et al. (2025); Bera and Mondal (2025b), but these do not give easy access to the underlying probability density and are therefore less suitable for MCMC sampling than normalizing flows.
@@
A promising approach is therefore to combine the strength of CV-enhanced sampling and generative models (normalizing flows) with a CV of ‘intermediate’ dimension (from a dozen to a few hundred degrees of freedom). Within this dimensionality, identifying a CV such that it resolves the metastability will be much easier than finding a one- to three-dimensional representation required by most classical CV-based approaches. Conversely, a normalizing flow in an ‘intermediate’ dimension can be trained to much higher accuracy than if one attempted to learn the distribution of the entire system. Since the overall goal is still sampling full configurations of the system, the question arises of how to employ such a proposal sampler in CV space to obtain an unbiased sampler for the full distribution. This has been addressed in multiple worksAthènes (2002); Nilmeier et al. (2011); Chen and Roux (2015a, b); Neal (2005), in particular under the names of “Hybrid Nonequilibrium Molecular Dynamics” (HNMD)Chen and Roux (2015a, b) and “Nonequilibrium candidate Monte Carlo” (NCMC) Nilmeier et al. (2011) in the context of molecular dynamics, but also in a Bayesian settingKaragiannis and Andrieu (2013). The central idea is always to drive the system from one value of the CV to another, a procedure referred to as steering in the following, and then accept or reject the resulting proposal in the full-space. A similar procedure has also been used in the simulation of glassy system with shear deformationsDas et al. (2022). In a previous articleSchönle et al. (2025), we described such an algorithm with driven dynamics in continuous time and identified various time-discretizations for which we proved reversibility with respect to the equilibrium measure. Notably, we also related this to the Jarzynski and the Jarzynski–Crooks equalities Jarzynski (1997); Crooks (1999), clarifying the algorithm design. In a parallel workTamagnone et al. (2024), one version of these algorithms was applied to a toy molecular system, using a normalizing flow as the CV sampler and including an adaptive protocol to train it alongside the MCMC procedure.
@@
The main part of the article divides into two sections: We explain all the theoretical aspects of the problem and the algorithm in Section˜II and go on to demonstrate its properties through several numerical examples in Section˜III. Within Section˜II, the sampling problem is introduced in Section˜II.1, followed by a detailed explanation of the algorithm in Section˜II.2. We state our main theoretical result on the reversibility of the algorithm in Section˜II.3 and conclude in Section˜II.4 with a normalized version of the algorithm and comments on the parameterization. In Section˜III, we analyze the performance of the algorithm by means of numerical simulations and present results on four different model systems of increasing complexity in Sections˜III.1, III.2, III.3 and III.4. We conclude the article and give an outlook on future directions in Section˜IV.
@@ II Sampling with a non-linear collective variable
Figure 1: Schematic representation of the algorithm for the example of a polymer, here with the end-to-end distance as the CV.
@@ II.1 Target distribution and collective variable
$\nu(\mathrm{d}q)=\mathcal{Z}_{q}^{-1}e^{-\beta V(q)}\mathrm{d}q$ (II.1)
over positions $q\in\mathbb{R}^{d}$ of dimension $d$, with the energy function $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the inverse temperature $\beta=(k_{B}T)^{-1}>0$ and the normalizing constant $\mathcal{Z}_{q}$. Let us assume that a collective variable (CV) $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\ell}$ with $\ell<d$ has been identified and we have a ‘good’ sampler $\rho$ (close to the target distribution in CV space) to propose moves in CV space. As mentioned in the introduction, the question of how to use these two ingredients to construct an unbiased MCMC sampler in the full space $\mathbb{R}^{d}$ has previously been addressed in literatureAthènes (2002); Nilmeier et al. (2011); Chen and Roux (2015a, b); Neal (2005), but it is not obvious how to employ these schemes if the map $\xi$ is non-linear. Notably, the geometry of the level sets $\Sigma(z)$ defined as
@@
must be taken into consideration. We present an algorithm which explicitly treats the case of a non-linear CV in the following section. Much of the notation throughout this paper is based on Refs. Lelièvre et al., 2010, 2012 and we refer the reader to those for an in-depth introduction on the mathematical concepts used here.
@@
$\mu(\mathrm{d}q\,\mathrm{d}p)=\mathcal{Z}^{-1}e^{-\beta H(q,p)}\mathrm{d}q\,\mathrm{d}p$ (II.3)
where $\mathcal{Z}$ is the normalizing constant, $H(q,p)=\frac{1}{2}p^{\top}M^{-1}p+V(q)$} is the Hamiltonian, and $M\in\mathbb{R}^{d\times d}$ is a symmetric positive definite mass matrix.
@@ II.2.1 Overall structure
The algorithm builds an MCMC chain $(Q_{n})_{n\geq 0}$ in position space by iterating over $n\in\mathbb{N}$ the following steps. Starting from a state $Q_{n}$ with collective variable value $Z_{n}:=\xi(Q_{n})$, generate $Q_{n+1}$ as follows:
@@ 2. Step 2:
    Generate a new $\widetilde{Q}_{n+1}\in\mathbb{R}^{d}$ that satisfies $\xi(\widetilde{Q}_{n+1})=\widetilde{Z}_{n+1}$ by driving the system in CV space from $Z_{n}$ to $\widetilde{Z}_{n+1}$. To this end,
    • (i)
      specify a schedule for the positions and velocities in CV space to fix the transition path (Section˜II.2.2),
    • (ii)
      initialize the position $q^{0}=Q_{n}$ and sample an initial auxiliary momentum $p^{0}$ (Section˜II.2.3),
    • (iii)
      use this to follow (Langevin) dynamics constrained in CV space by the schedule (Section˜II.2.4),
    all of these steps to be specified below. These dynamics yield a final state $\widetilde{Q}_{n+1}$ and a quantity $\mathcal{W}_{n+1}$ that is identified as a work. The final momentum is discarded.
@@ 3. Step 3:
The three steps of the algorithm are schematically illustrated in Fig.˜1.
In principle, any probability density $\rho(z,z^{\prime})$ can be chosen to generate the proposal moves in CV space in Step 1, since the accept/reject criterion in Step 3 always ensures unbiasedness, as will be proven below. This choice will, however, affect the acceptance probability and thus the performance of the algorithm. In our numerical examples, we always use an independent sampler in the CV space, i.e. $\rho(z,z^{\prime})$ is a function of $z^{\prime}$ only. A natural choice is (an approximation of) the marginal density $\rho(z^{\prime})$ of the target measure in CV space – that is the image measure of the target distribution (II.1) under the CV map $\xi$. For a CV of intermediate dimension, such a proposal sampler could for example be learned by a normalizing flow.
We explain the various ingredients required for Step 2 in the next three subsections. The idea to sample a system by following some dynamics followed by an accept/reject step that is governed by a work is closely related to the Jarzynski and the Jarzynski–Crooks equalities Jarzynski (1997); Crooks (1999).
@@ II.2.2 Schedule
To specify the constrained dynamics under which the system evolves from the value of the CV at the previous iteration $Z_{n}$ to the newly proposed value $\widetilde{Z}_{n+1}$, choose first the number of intermediate steps $K_{T}$ associated with this jump. With a fixed time discretization $\Delta t$, this sets the physical transition time $T=K_{T}\Delta t$. In practice, we will choose $K_{T}$ as a function of the Euclidean distance $|\widetilde{Z}_{n+1}-Z_{n}|$ of the proposed jump.
@@
for all $k\in\{0,\ldots,K_{T}\}$. In practice, it is advisable for numerical reasons to choose the velocity schedule in accordance with the position schedule (see Remark˜B.4 in the appendix). Note, however, that the requirement (II.8) precludes the simplest choice for the velocity of $v_{z}(t_{k})=(z(t_{k+1})-z(t_{k}))/\Delta t$. A simple choice that admits a continuous-in-time limit when $\Delta t$ goes to zero is instead a schedule that connects the endpoints on a straight line with a continuously differentiable schedule function $f:[0,1]\rightarrow[0,1]$:
@@
$v^{(Z,\widetilde{Z})}_{z}(t_{k})$ $=(\widetilde{Z}-Z)\frac{f^{\prime}(k/K_{T})}{K_{T}\Delta t}.$ (II.10)
The conditions on the schedule translate to conditions on $f$ as $f(0)=0$, $f(1)=1$ and $f^{\prime}(0)=f^{\prime}(1)=0$, which are fulfilled for example by
@@
This is one simple choice to parameterize the schedule, but many others are possible as long as the requirements stated above are fulfilled.
The choice of transition path in CV space as well as the time schedule to follow it determines how far from equilibrium the dynamics will take the system, which affects the acceptance ratio and thus the overall numerical performance (theoretically, the highest possible acceptance will be achieved by driving infinitely slowly, but at the expense of longer computational times to integrate the steered dynamics). In our examples, we stick to the simple prescription explained above, where the transition path is a straight line in CV space, and only optimize the overall speed with which that path is followed. See Remark˜B.4 in Section˜B.2.1 for further comments on the velocity schedule. The assumptions (II.6) and (II.8) play a crucial role in the proof of the reversibility of the Markov chain $(Q_{n})_{n\geq 0}$ and we demonstrate in Section˜H.2 how their violation can introduce a bias. Note that for some choices of CV, the condition (II.6) can be somewhat relaxed, see Remark˜II.2.
Remark II.1.
@@ II.2.3 Initialization of the momentum
In each iteration of the algorithm, a fresh momentum is sampled in Step 2 to initialize the constrained dynamics. To agree with condition (II.6), this momentum must correspond to zero velocity in $\xi$-direction (i.e. in CV space) and match with the underlying equilibrium canonical distribution orthogonally to $\xi$. We explain here how to achieve this. For a given position $q$ and momentum $p$, the velocity $v_{\xi}$ is defined as
@@
where $\nabla\xi(q)=\left(\nabla\xi_{1}(q),\ldots,\nabla\xi_{\ell}(q)\right)\in\mathbb{R}^{d\times\ell}$. The cotangent space of zero velocity associated with a position $q\in\Sigma(z)$ is defined as
@@
We therefore initialize the momentum $p^{0}$ by sampling an unconstrained momentum $\widetilde{p}^{0}$ from a centered normal distribution with covariance matrix $\frac{1}{\beta}M$ and projecting it onto the subspace of zero velocity:
@@
II.2.4 Constrained dynamics and work
We now specify how the system is driven under constrained dynamics to follow the chosen CV and velocity schedules. Special care needs to be taken here regarding the geometry of the level sets $\Sigma(z)$. To account for the change in phase space volume due to the rigid constraints, we need to follow the dynamics with a modified potential energy $\widetilde{V}(q)=V(q)+V_{\mathrm{fix}}(q)$, with the so-called Fixman termFixman (1978).
$V_{\mathrm{fix}}(q)=\frac{1}{2\beta}\log\left(\det G_{M}(q)\right),$ (II.16)
(see also the reversibility proof in Section˜B.3 and Remarks˜B.2 and B.10 for the technical reason why it is necessary). We numerically demonstrate the necessity of this modification in Section˜H.2. Note that the tilde on $\widetilde{V}$ and $\widetilde{H}$ indicates here the modification by the Fixman term, while we also use it on $\widetilde{Q}_{n+1}$ and $\widetilde{Z}_{n+1}$ to designate proposal states in the MCMC chain.
Starting from $(q^{0},p^{0})$, a path $\left((q^{k},p^{k})_{0\leq k\leq K_{T}}\right)$ is generated by discretizing in time a constrained Langevin dynamics, with a splitting scheme of the form (midpoint Euler)-(Verlet)-(midpoint Euler) (see for example Ref. Lelièvre et al., 2012). The first and last parts (the thermostat parts) insert fluctuations and the middle part realizes Hamiltonian dynamics. For the latter, we use a RATTLE discretization scheme with constraints, slightly different to the standard version since we consider constraints which evolve in time (see also Remark˜B.4 in the Appendix for further discussion). For the sake of readability, we drop the superscript indicating the endpoints of the schedules $z(t)$ and $v_{z}(t)$ in this subsection. The scheme writes as follows: For $0\leq k\leq K_{T}-1$, update the positions and momenta as